TIME DOMAIN THz SPECTROSCOPY OF INTRACOMPLEX RYDBERG-RYDBERG TRANSITIONS IN CaF

\underline{VLADIMIR S. PETROVIC,} EMILY E. FENN, AND ROBERT W. FIELD, Massachusetts Institute of Technology, Chemistry Department, 77 Massachusetts Avenue, Cambridge, MA 02139.

Exceptionally large transition dipole moments among Rydberg states lend themselves to simplifying the analysis of the strongly mixed regions. We used the Stark effect in a double resonance experiment and are developing a Time Domain THz Spectrometer for a three photon experiment, to benefit from this property. Progress in the experiments, as well as predicted spectra, will be presented.